# Smaller and More Flexible Cuckoo Filters

### v1

@@ Appendix C Comparison of load thresholds with achievable loads
Figure 6 (left) shows the achievable loads for different maximum random walk lengths. As expected, the achievable load increases for longer random walk lengths and converges to the load threshold if enough steps are allowed for the random walk during insertion. We conclude that 10 000 to 20 000 steps are enough to achieve loads close to the thresholds. Note that large numbers of steps are only taken when the filter is almost at capacity; as long as it is sufficiently empty, insertion happens in a single step (or with a few steps), independently of the maximum walk length.
@@
For a fixed random walk length, the achievable load should be independent of the FPR and the relative filter size. In practice, the achievable load appears independent of the FPR, except for small $k$ with e.g. an FPR of $2^{-4}$, for which the achieved load is lower than for higher $k$ (Figure 6 right). A possible explanation is that having only $152^{4}-1=$ 15 distinct fingerprints reduces the randomness for the second bucket or window. In addition, the achievable load decreases slightly for larger filters (with constant overhead factor, i.e., both $n$ and the filter size are at constant ratio), which may also be explained by the problem of mapping the small space of fingerprints to the much larger number of buckets or windows.

### v2

@@ Appendix B Comparison of load thresholds with achievable loads
Figure 6 (left) shows the achievable loads for different maximum random walk lengths. As expected, the achievable load increases for longer random walk lengths and converges to the load threshold if enough steps are allowed for the random walk during insertion. We conclude that 10 000 steps are enough to achieve loads close to the thresholds with high probability. Note that large numbers of steps are only taken when the filter is almost at capacity; as long as it is sufficiently empty, insertion happens in a single step (or with a few steps), independently of the maximum walk length.
@@
For a fixed random walk length, the achievable load should be independent of the FPR and the relative filter size. In practice, the achievable load appears independent of the FPR, except for small $k$ with e.g. an FPR of $2^{-4}$, for which the achieved load is lower than for higher $k$ (Figure 6 right). A possible explanation is that having only $2^{4}-1=15$ distinct fingerprints reduces the randomness for the second bucket or window. In addition, the achievable load decreases slightly for larger filters (with constant overhead factor, i.e., both $n$ and the filter size are at constant ratio), which may also be explained by the problem of mapping the small space of fingerprints to the much larger number of buckets or windows.
Appendix C Distribution of random walk lengths
Long random walks cause many cache misses and hence lead to slow insertion times. Therefore, short random walks are preferable. This can easily be achieved by increasing the space overhead of the filter, since the costly insertions happen when the filter is almost filled to its maximum capacity (see Figure 6). However, in many applications small filters are preferable, in particular, if more queries than insertions are performed.
Figure 7: Distribution of random walk lengths when inserting $10^{8}$ elements into different Cuckoo filter types for a fixed FPR of $2^{-10}$ and a filter size that ensures a load of at most $98\%$ of the theoretical load threshold.
Figure 7 shows the random walk length distribution of bucketed and windowed Cuckoo filters allowing a maximum random walk length of $10^{4}$. Most keys are directly inserted into a free slot and the filter performs a long random walk only for a very small fraction of keys; note the logarithmic scale of the y-axis. Choosing a maximum load such that the filter is only filled up to $98\%$ of its maximum capacity ensures that the actual maximum random walk length is much lower than the allowed maximum random walk length of $10^{4}$, with $(2,4)$ Cuckoo filters having shorter random walks compared to $(2,2)$ Cuckoo filters.
Appendix D Comparison of C++ and Python (with numba) implementations
To rule out that our Cuckoo filter implementation is slower or faster due to the use of just-in-time compiled Python, we compared the runtime of the original implementation of the Cuckoo filter ($(2,4)$-bucketed Cuckoo filter, computing the alternative bucket with an XOR operation, implemented in C++ [12]) with our implementation of the the same (2,4)-bucketed Cuckoo filter. Both implementations are as similar as possible, e.g., both implementations use the same trick for bit-parallel lookups, but for example, different hash functions.
Table 4: Throughput in million keys per second, single threaded; higher is better. We compare the original Cuckoo filter implementation written in C++ with an implementation in just-in-time compiled Python using numba. All filters have $2^{30}$ slots and a load of $95\%$.
FPR $2^{-5}$ (8-bit slots) FPR $2^{-13}$ (16-bit slots)
insert lookup insert lookup
C++ 7.305 61.939 7.371 60.517
numba 8.938 70.688 9.104 55.013
Table 4 shows the throughput for both implementations for Cuckoo filters configured to have a load factor of $95\%$. The throughput is averaged over three repeated runs. We measure lookup time by querying $50\%$ contained and $50\%$ not contained keys. Since both implementations have similar running times, in particular for lookups, the numba implementation does not affect our reasoning on an algorithm engineering level when comparing our novel Cuckoo filters with other state-of-the-art probabilistic filters.